OpenStudy (anonymous):

help!!! solve the differential eqn: x''-2x'+x=e^(2t)+t given x(0)=0 and x'(0)=4

OpenStudy (turingtest):

you can use variation of parameters or undetermined coefficients which do you prefer?

OpenStudy (turingtest):

either way you should first get the complimentary solution, i.e. solve\[x''-2x'+x=0\]

OpenStudy (anonymous):

undetermined co-efs....(f-1)(f-1)=0 x=Ae^x+Bxe^x

OpenStudy (turingtest):

ok, the we have the particular given by\[y_p=-y_1\int\frac{y_2g(t)}Wdt+y_2\int\frac{y_1g(t)}Wdt\]where g(t)=e^(2t)+t, and W is the Wronskian. Do you have the Wronskian yet?

OpenStudy (turingtest):

then*

OpenStudy (anonymous):

noooo, we haven't done that yet... wat does W indicate?

OpenStudy (turingtest):

\[W(y_1,y_2)=\left|\begin{matrix}y_1&y_2\\y_1'&y_2'\end{matrix}\right|\]

OpenStudy (anonymous):

(O__o)... oh my!!!

OpenStudy (turingtest):

erm, if you don't know at all what I'm saying then I think you need to review variation of parameters the above is a determinant of the matrix. do you remember determinants for 2x2 matrices?

OpenStudy (turingtest):

Oh I'm sorry you wanted undetermined coefficients huh...

OpenStudy (anonymous):

ow yes... that cheap stuff... yes... in a 2x2 matrix with [a b,c d]... becomes ad-bc

OpenStudy (turingtest):

my mistake I was doing variation of parameters so for undetermined coefficients, what is your guess for the particular going to be?

OpenStudy (anonymous):

i tried 4 guesses... stil got contradictions... :( i tried x= Ae^(2t)+Ax+t

OpenStudy (turingtest):

you got x and t mixed together, is that what you tried to use?

OpenStudy (anonymous):

yes :(

OpenStudy (turingtest):

yp=Ae^(2t)+Bt+C

OpenStudy (turingtest):

the guess for the e^(2t) part is Ae^(2t) the guess for the t is Bt+C add them together

OpenStudy (anonymous):

0=A+B+C for x 4=2A+B for x'

OpenStudy (turingtest):

did you plug in the initial conditions or something?

OpenStudy (anonymous):

wait... after substituition i'm stuck here 1-B+Ae^(2t)+C=e^(2t)+t

OpenStudy (turingtest):

I don't see how you got 1-B

OpenStudy (turingtest):

x''=4Ae^(2t) x'=2Ae^(2t)+B x=Ae^(2t)+Bt+C

OpenStudy (anonymous):

after substitution yp=Ae^(2t)+Bt+C y'=2Ae^(2t)+B y''=4Ae^(2t)+1

OpenStudy (turingtest):

B is a constant

OpenStudy (anonymous):

when u differentiate B, becomes 0 or 1?

OpenStudy (turingtest):

lone constants differentiate to zero

OpenStudy (anonymous):

owwww, ok, let me try again

OpenStudy (anonymous):

i got A=1 B=1 C=1

OpenStudy (turingtest):

C is not 1, but the others are

OpenStudy (anonymous):

When i equated coeffic.... i got (-B+C)=0 Then -B=-C... C and B hav the same values

OpenStudy (turingtest):

I think you dropped a 2 from the -x' part, look again

OpenStudy (anonymous):

still getting the same answer... 4Ae^(2t)-4Ae^(2t)-2B+Ae^(2t)+Bt+C=e^(2t)+t

OpenStudy (anonymous):

that's my substitution

OpenStudy (turingtest):

this is right

OpenStudy (turingtest):

so it simplifies to what?

OpenStudy (anonymous):

Bt=t -2B+B+C=0 Ae^(2t)=e^(2t)

OpenStudy (turingtest):

what is the Bt=t thing?

OpenStudy (turingtest):

Ae^(2t)+Bt-2B+C=e^(2t)+t is what I was going for

OpenStudy (anonymous):

-2B?

OpenStudy (turingtest):

set the coefficients equal what is the coefficient of the e^(2t) term on the left?

OpenStudy (anonymous):

it's A

OpenStudy (turingtest):

and on the right?

OpenStudy (anonymous):

1

OpenStudy (turingtest):

right.so A=1 now what is the coefficent of t on the left and rigth? what is B then ?

OpenStudy (anonymous):

B=1

OpenStudy (turingtest):

good and what is the constant term on each side?

OpenStudy (anonymous):

-2B+C=0

OpenStudy (turingtest):

so what is C ?

OpenStudy (anonymous):

C=2

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

wow... so my soln becomes yp=e^(2t)+t+2

OpenStudy (turingtest):

that's the particular, yep

OpenStudy (anonymous):

y=Ae^(x)+bxe^(x)+e^(2t)+t+2??

OpenStudy (turingtest):

yes, now you can apply the initial conditions

OpenStudy (turingtest):

and keep your x's and t's straight...

OpenStudy (turingtest):

x(t)=c1e^(t)+c2te^(t)+e^(2t)+t+2

OpenStudy (anonymous):

let me try it out

OpenStudy (anonymous):

x(0)=0 i leave the t as it is in my substitution?

OpenStudy (turingtest):

x=0 and t=0, so you have to plug in 0 for t

OpenStudy (anonymous):

i'm getting A=-3... B=0

OpenStudy (turingtest):

A=-3 yes, B=0 no

OpenStudy (turingtest):

you can't find B without the other condition. Did you apply iot yet?

OpenStudy (turingtest):

it*

OpenStudy (anonymous):

yes... x'(0)=4 x'=Ae^(2t)+Bte^(2t)+e^(2t)

OpenStudy (turingtest):

that is not right for x'

OpenStudy (turingtest):

and those first two e^(2t)'s should be e^t's

OpenStudy (anonymous):

ow yes... i'm kinda confused :(

OpenStudy (turingtest):

\[x(t)=Ae^{t}+Bte^{t}+e^{2t}+t+2\]\[x'(t)=Ae^t+B(t+1)e^t+2e^{2t}+1\]

OpenStudy (anonymous):

u really fast... u on matlab or mathematica? i got B=6

OpenStudy (turingtest):

no computer help, I just know these things and already solved it it is not B=6, maybe you got that from my typo though

OpenStudy (anonymous):

not 6? :(

OpenStudy (turingtest):

\[x'(0)=A+B+2+1=4\]

OpenStudy (anonymous):

B=4?? :) yes?

OpenStudy (turingtest):

yes :) and that's basically the end of the problem

OpenStudy (anonymous):

not really :)

OpenStudy (turingtest):

add the complimentary and the particular the end

OpenStudy (anonymous):

are u tired of helping me? :(... can u atleast check my last step... they asked me to find x(1)

OpenStudy (turingtest):

well, what is x(t) ?

OpenStudy (anonymous):

x(1)= 13.107

OpenStudy (anonymous):

final answer :)

OpenStudy (turingtest):

yep, that's right :)

OpenStudy (anonymous):

thank u... its midnyt here in south africa... got a test at 2pm... thank u so much... wil post another question if i do get stuck sumwher

OpenStudy (turingtest):

You're welcome. I'll be happy to help if I'm available. See ya!

OpenStudy (anonymous):

:)